OpenStudy (anonymous):

please Help !!! Fan And Medal !! Your friend hands you a graph of the engine power statistics of a race car. He says, “I know this graph is f(x) = –3(x + 4)6 – 8 but I can’t remember how it is related to the graph of x6.” Explain to your friend how the graph f(x) is a translation of the graph x6.

OpenStudy (anonymous):

Can you Please Help Me ?? This Is My Final Question And Im Stuck !!

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

@phi ?????

OpenStudy (phi):

\[ f(x) = –3(x + 4)^6 – 8 \] compared to \( f(x)= x^6 \) if we put -3 in front of the x^6 , the negative means we "flip it upside down" the 3 causes the curve to be "steeper" (the answer is bigger than if we did not multiply) replacing x with (x+4) causes the "zero" to be at -4 (in other words, when x=4 in the new version (x+4), we get (-4+4) = 0 that means the x=0 in the original matches up with x=-4 in other words, we translated LEFT 4 steps adding -8 to this –3(x + 4)^6 – 8 makes f(x) smaller by -8 as compared to –3(x + 4)^6 in other words we translated DOWN by 8 steps